Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 414     Entry time: 25 Oct 2023, 15:10
Author: Sergii, Teguh 
Group: SCS 
Subject: Morning Shift Summary, 24.10.2023 

Morning Shift Summary, 24.10.2023

Achievements:

  1. Re-aligned the beamline after maintenance
  2. Found time zero using SiN at 300 K
  3. Change the He dewar and cooled down back to 20 K
  4. Adjusted the AFS laser parameters to recover the proper experiment settings
  5. Started taking tr-RIXS with several time delay points

Problems:

  1. The AFS laser power is smaller than last week

Next shift plan:

  1. Continue taking tr-RIXS data
ELOG V3.1.4-7c3fd00